Marcus Amy and Colin are collecting quarters for a school fundraiser. Marcus has 12 more than Amy Colin has three times more than Amy all together they have 107. how many quarters does Marcus have and how many quarters does Colin have
Answer by stanbon(75887)

Marcus Amy and Colin are collecting quarters for a school fundraiser. Marcus has 12 more than Amy Colin has three times more than Amy all together they have 107. how many quarters does Marcus have and how many quarters does Colin have
---------------------------
m + a + c = 107
m = a+12
c = 3a
-----------------
Solve for "a":
a+12 + a + 3a = 107
5a = 95
---------
a = 19 (# Amy has)
m = a+12 = 31 (# Marcus has)
c = 3a = 57 (# Colin has)
----------------------------
Cheers,
Stan H.
--------------------------